OpenStudy (jkbo):

write the equation of an ellipse with minor axis of length 8 and vertices at (-9,3) and (7,3)

OpenStudy (mertsj):

The minor axis = 8 tells you that 2b=8

OpenStudy (mertsj):

Can you find the distance from (-9,3) to (7,3)? That distance is the length of the major axis which is 2a

OpenStudy (jkbo):

its 17 units away

OpenStudy (mertsj):

From (-9,3) to (7,3) is 16

OpenStudy (mertsj):

The center is the midpoint of the major axis.

OpenStudy (jkbo):

so 16 is the midpoint?

OpenStudy (mertsj):

The length of the major axis is 16

OpenStudy (mertsj):

Do you know how to find a midpoint?

OpenStudy (jkbo):

no not at all

OpenStudy (mertsj):

Add the x coordinates and divide by 2. Add the y coordinates and divide by 2. That will give you the midpoint.

OpenStudy (mertsj):

\[x=\frac{x _{1}+x _{2}}{2}\]

OpenStudy (mertsj):

\[y=\frac{y _{1}+y _{2}}{2}\]

OpenStudy (jkbo):

I got -1,3

OpenStudy (mertsj):

Yes. That is the center.

OpenStudy (mertsj):

Now you just need to plug into the equation of an ellipse:

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]

OpenStudy (mertsj):

(h,k) is the center.

OpenStudy (jkbo):

what would x and y be?

OpenStudy (jkbo):

@Mertsj

OpenStudy (mertsj):

x and y remain in the equation.

OpenStudy (jkbo):

ok does a^2 and b^2 stay the same or no?

OpenStudy (mertsj):

Replace a and b with the values you calculated.

OpenStudy (jkbo):

-1,3 again?

OpenStudy (mertsj):

2b=8

OpenStudy (mertsj):

Can you find the distance from (-9,3) to (7,3)? That distance is the length of the major axis which is 2a

OpenStudy (jkbo):

that 16

OpenStudy (jkbo):

*thats

OpenStudy (mertsj):

Yes. 2a = 16

OpenStudy (jkbo):

b^2 is 8 and a^2 is 16

OpenStudy (mertsj):

2b=8 Solve for b

OpenStudy (mertsj):

Yes. 2a = 16 Solve for a

OpenStudy (jkbo):

its 4

OpenStudy (mertsj):

If b = 4, what is b^2?

OpenStudy (jkbo):

I'm not really sure

OpenStudy (mertsj):

What class are you taking?

OpenStudy (jkbo):

alg. 2 I just dont know this chapter well

OpenStudy (mertsj):

How is it you don't know how to raise a number to the second power?

OpenStudy (jkbo):

o well 2 to the second power is 4 if that what your asking

OpenStudy (mertsj):

You said that b = 4 and that is correct. I then asked, "what is b^2?" You said you didn't know.

OpenStudy (jkbo):

I thought you want the square root of 16 thats why or what can go into 8 and 2

OpenStudy (mertsj):

What is a?

OpenStudy (jkbo):

I pretty sure its 4 because of 16

OpenStudy (mertsj):

2a = 16 Solve for a

OpenStudy (jkbo):

I'm not sure what it is I confused

OpenStudy (jkbo):

i'm *

OpenStudy (mertsj):

So you are telling me that you are in Algebra II, studying ellipses and you cannot solve this equation:

OpenStudy (mertsj):